OpenStudy (anonymous):

A cube has a surface area of 6x. What is the volume of the cube in terms of x?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

v=6x*x v=6x^2

OpenStudy (dumbcow):

sa = 6a, where a is area of each square face a = s^2, where s is length of each side volume = s^3 6x = 6a x=a x=s^2 sqrt(x) = s s^3 = (sqrt(x))^3 volume = x^3/2 = xsqrt(x)

OpenStudy (anonymous):

the area of each face is x so each side = sqrt x so vol = (sqrtx)^3 = x^(3/2)